# Stable Score Distillation for High-Quality 3D Generation

### 14 Dec 2023 v1

@@ Abstract
Score Distillation Sampling (SDS) has exhibited remarkable performance in conditional 3D content generation. However, a comprehensive understanding of the SDS formulation is still lacking, hindering the development of 3D generation. In this work, we present an interpretation of SDS as a combination of three functional components: mode-disengaging, mode-seeking and variance-reducing terms, and analyze the properties of each. We show that problems such as over-smoothness and color-saturation result from the intrinsic deficiency of the supervision terms and reveal that the variance-reducing term introduced by SDS is sub-optimal. Additionally, we shed light on the adoption of large Classifier-Free Guidance (CFG) scale for 3D generation. Based on the analysis, we propose a simple yet effective approach named Stable Score Distillation (SSD) which strategically orchestrates each term for high-quality 3D generation. Extensive experiments validate the efficacy of our approach, demonstrating its ability to generate high-fidelity 3D content without succumbing to issues such as over-smoothness and over-saturation, even under low CFG conditions with the most challenging NeRF representation.
Figure 1: 3D Gaussian Splatting generation from text prompts.
@@ 1 Introduction
3D generative modelling is inherently more complex than 2D modelling, requiring meticulous attention to view-consistent fine geometry and texture. Despite the increased intricacy, 3D data is not as readily available as its 2D image counterparts, which have propelled recent advances in text-to-image generation [33, 35, 34]. Even with recent efforts of Objaverse [7], the uncurated 3D data only amounts to 10 million instances, in sharp contrast to the vast 5 billion image-text pairs available [36]. As a result, utilizing 2D supervision for 3D generation has emerged as a prominent research direction for text-to-3D generation. Notably, Score Distillation Sampling (SDS) [30, 44] is proposed to optimize a 3D representation so that its renderings at arbitrary views are highly plausible as evaluated by a pre-trained text-to-image model, without requiring any 3D data. As presented in DreamFusion [30], SDS enables the production of intriguing 3D models (NeRFs [25, 2, 26]) given arbitrary text prompts, but the results tend to be over-smooth and over-saturated. Subsequent works build upon SDS and enhance the generation quality with improvements in training practices. This is most effectively achieved by adopting higher resolution training, utilizing larger batch sizes, and implementing a coarse-to-fine generation approach with mesh optimization for sharper generation details [21, 5, 45]. While SDS remains fundamental and ubiquitously used for 3D generation, its theoretical understanding remains obscured and requires further exploration.
In this paper, we provide a thorough understanding of SDS and introduce our revised formulation, termed Stable Score Distillation (SSD). The proposed method addresses the previously overlooked challenges of over-smoothness and color-saturation in 3D generation. Specifically, we decompose the diffusion-guided noise residual term in SDS (referred to as the SDS estimator) into mode-seeking, mode-disengaging and variance-reducing components. We then delve into their mathematical properties and their impacts on optimization-based 3D generation. With our analysis, we show that the mode-seeking term is attracted by transient modes and inevitably causes the over-smoothing problem. And the mode-disengaging term drives the rendering of the 3D representation away from image modes specified by pre-trained 2D diffusion models when diffusion timestep $t$ is small, causing color-saturation. Additionally, we uncover the inefficiency of the variance-reduction scheme in SDS resulting from the scale and direction mismatch between the variance-reducing and the mode-seeking terms. Based on the insights, we propose a simple yet effective approach that elegantly avoids the intrinsic problems of each term with a strategic combination. Our method significantly alleviates the issues of over-smoothness and color-saturation, with low Classifier-Free Guidance (CFG). Furthermore, our interpretation of the training signals establishes meaningful connections with common observations and challenges in optimization-based 3D generation, revealing opportunities for further advancements.
Our main contributions are summarized as follows:
@@
  • We offer a comprehensive understanding of Score Distillation Sampling (SDS) by interpreting the noise residual term (referred to as the SDS estimator) as a combination of three functional components: mode-disengaging, mode-seeking and variance-reducing terms. We conduct thorough evaluation of their distinctive mathematical properties, intrinsic limitations and their impacts on optimization-based 3D generation.
  • We identify that the over-smoothness and color-saturation problems in 3D generation arise from intrinsic deficiency of the mode-disengaging and mode-seeking terms. Additionally, we show that the variance reduction scheme of SDS is inefficient as it overlooks the scale and direction mismatch between the mode-seeking and variance-reducing terms. Through the analysis, we acknowledge the main training signal of 3D generation to be the mode-disengaging term.
@@
  • Building on our analysis, we propose Stable Score Distillation (SSD), which utilizes a simple yet effective term-balancing scheme that significantly mitigates the fundamental deficiencies of SDS, particularly addressing issues such as over-smoothness and color-saturation.
@@
  • We establish meaningful connections between our analytical findings and prevalent observations and practices in optimization-based 3D generation, such as the adoption of a large CFG scale and the generation of over-smoothed and over-saturated results. These connections provide valuable perspectives for enhancing the optimization process of 3D generation.
@@
  • Through extensive experiments, we demonstrate that our proposed method significantly outperforms SDS via straightforward reformulation of training signals. Our approach is capable of high-quality 3D generation with more vibrant colors and finer details.
Figure 2: 3D NeRF generation from text prompts.
@@ 2 Related Works
Recently, DreamFusion [30] and SJC [44] propose to generate 3D content by optimizing a differentiable 3D representation so that its renderings at arbitrary viewpoints are deemed plausible by 2D diffusion priors. Such methods are commonly referred to as the optimization-based 3D generation. Subsequent works enhance the generation quality by utilizing 3D-aware diffusion models for image-to-3D generation [42, 32, 18, 24]; adopting a coarse-to-fine generation approach with improved engineering practices such as higher-resolution training and mesh optimization [32, 5]; introducing additional generation priors [14, 3, 15, 20, 50]; or exploring timestep scheduling [13, 51]. ProlificDreamer [45], along with recent works of NFSD [16] and CSD [49] concurrent to our study provide new perspectives on SDS. However, there remains a gap in a comprehensive understanding of the SDS formulation, which is the focus of this work.
@@ 3.4 Score Distillation Sampling (SDS)
where $w(t)$ re-scales the weights of gradients according to $t$, and $p_{\phi}(\boldsymbol{x}_{t};y,t)$ is the distribution of $\boldsymbol{x}_{t}$ implicitly represented by the score function $\hat{\boldsymbol{\epsilon}_{\phi}}$. DreamFusion [30] introduces the $\boldsymbol{\epsilon}$ term for variance reduction. We call the resultant noise residual term highlighted in red the SDS estimator for short.
@@ 4 Method
We start with problem setting in Sec. 4.1 and decompose the SDS estimator to three functional terms as shown in Eq. 4: namely the mode-disengaging term (denoted as $h$ for its homogeneous form), the mode-seeking term and the variance-reducing term. In Sec. 4.2, Sec. 4.3, and Sec. 4.4 we analyze the mathematical properties, numerical characteristics and intrinsic limitations of each term, respectively. Based on our analysis, we propose Stable Score Distillation (SSD) in Sec. 4.5 which efficiently utilizes the distinctive properties of each term to augment each other for higher-quality 3D content generation.
@@ 4.1 Problem Setting
Figure 3: Evaluating the learning behaviour of $\theta$ supervised by different estimators with an illustrative 2D example. In this toy example $\theta=\boldsymbol{x}\in\mathbb{R}^{2}$, and $\mathbb{P}(\boldsymbol{x})$ is a mixture of Gaussian distributions. Mode-Disengaging trajectory: the learning trajectory of $\boldsymbol{x}$ by employing the mode-disengaging term only in Eq. 3, where Mode-Seeking trajectory and SSD trajectory are defined similarly. The purple trajectory describes how $\theta$ evolves when initialized at the trapping point and supervised by mode-disengaging term. We present the detailed experiment analysis in Sec. 5.2.
@@ 4.2.1 Mathematical Properties
Substituting Eq. 2 into the formulation of $h$, we obtain $h=-\sigma_{t}(\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)-\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};t))$. When optimizing a 3D representation $\theta$, the sub-terms of $h$, namely $-\sigma_{t}\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$ and $\sigma_{t}\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,\emptyset)$, respectively drives $\theta$ such that $\mathbb{E}_{\boldsymbol{\epsilon}}[\boldsymbol{x}_{t}]=\alpha_{t}\boldsymbol{x}$ will: a) approach the conditional modes of $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$, and b) distance the unconditional modes of $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};t)$ which are independent of the conditional prompt (for this reason we call $h$ mode-disengaging). We emphasize property b) to be critical, as it guides $\theta$ to avoid the over-smoothing problem caused by trapping points and transient modes, a concern encountered by the mode-seeking term (see Sec. 4.3 for definition and analysis). In Fig. 3 (left), we illustrate that the mode-disengaging term has the capability to navigate out of the trapping point of the mode-seeking term induced from its transient modes, thanks to the disengaging force that avoids unconditional modes. This capability is crucial for effective 3D generation. Considering that in Eq. 4, only the mode-disengaging and mode-seeking terms provide supervision signals (note that $\mathbb{E}_{\boldsymbol{\epsilon}}[\boldsymbol{\epsilon}]=\boldsymbol{0}$), our analysis positions $h$ as the main supervision signal in 3D generation. Our analysis provides theoretical explanation to a concurrent work, CSD [49], which also claims $h$ to be the main supervision signal for text-to-3D generation.
@@ 4.2.2 Intrinsic Limitation
Based on previous discovery, one may be tempted to employ $h$ as the sole supervision signal in Eq. 3 for 3D generation, akin to CSD [49]. However, our experiments reveal a persistent issue of over-saturation in even 2D content generation when relying solely on $h$ (SSD, M=0 in Fig. 4). To delve deeper into this issue, we conduct experiments with varying values of $t$ and identify its particular association with small $t$, e.g. $t<200$ (see App. Fig. 17). Note that as $t\rightarrow 0$, $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)\rightarrow\mathbb{P}_{\phi}(\boldsymbol{x};y,t)$ and $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};\emptyset,t)\rightarrow\mathbb{P}_{\phi}(\boldsymbol{x};\emptyset,t)$. As the term $h$ seeks to maximize $\frac{\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)}{\mathbb{P}_{\phi}(\boldsymbol{x}_{t};\emptyset,t)}$, when $t\rightarrow 0$, the learning target of $h$ gradually approaches maximizing $\frac{\mathbb{P}_{\phi}(\boldsymbol{x};y)}{\mathbb{P}_{\phi}(\boldsymbol{x})}$. However, the regions where this ratio attains the maximal value are often not around any modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y)$, where $\mathbb{P}_{\phi}(\boldsymbol{x})$ is also high according to the mode consistency hypothesis. Worse still, they tend to be not around any singular modes of $\mathbb{P}_{\phi}(\boldsymbol{x})$, where $\mathbb{P}_{\phi}(\boldsymbol{x})$ is high and $\mathbb{P}_{\phi}(\boldsymbol{x};y)$ is low. In other words, when $t\rightarrow 0$, the mode-disengaging term discourages $g(\theta,c)$ from converging to any mode of a natural image distribution $\mathbb{P}_{\phi}(\boldsymbol{x})$, let alone to modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y)$, violating the necessary condition for 3D generation as presented in Sec. 4.1. Therefore, special treatments are necessary to enable the 3D representation to converge to modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y)$ when $t$ is small. We illustrate this problem in Fig. 3 (left), where the mode-disengaging trajectory converges to a point far from modes of $\mathbb{P}(\boldsymbol{x};y)$.
@@ 4.3.1 Numerical Properties
Scale. Recall $||\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)||=\sigma_{t}\cdot||\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)||$. Assuming $||\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)||$ is bounded, one should expect $||\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)||$ to increase as $t$ increases. We validate this assumption numerically in App. Fig. 19.
Direction. As $t\rightarrow T$, $\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$ gradually transforms from the score of conditional distribution of natural images, i.e., $\nabla_{\boldsymbol{x}}log\mathbb{P}_{\phi}(\boldsymbol{x};y)$, to $\nabla_{\boldsymbol{x}_{T}}\mathcal{N}(\boldsymbol{x}_{T};\boldsymbol{0},\boldsymbol{I})$, which are independent of $\boldsymbol{\epsilon}$ and collinear to $\boldsymbol{\epsilon}$ respectively (proved in App. Sec. 12). Therefore the linear correlation between $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)$ and $\boldsymbol{\epsilon}$ as defined in [31] is expected to increase as $t$ increases. We validate this assumption in App. Fig. 21.

### 07 Feb 2024 v2

@@ Abstract
Although Score Distillation Sampling (SDS) has exhibited remarkable performance in conditional 3D content generation, a comprehensive understanding of its formulation is still lacking, hindering the development of 3D generation. In this work, we decompose SDS as a combination of three functional components, namely mode-seeking, mode-disengaging and variance-reducing terms, analyzing the properties of each. We show that problems such as over-smoothness and implausibility result from the intrinsic deficiency of the first two terms and propose a more advanced variance-reducing term than that introduced by SDS. Based on the analysis, we propose a simple yet effective approach named Stable Score Distillation (SSD) which strategically orchestrates each term for high-quality 3D generation and can be readily incorporated to various 3D generation frameworks and 3D representations. Extensive experiments validate the efficacy of our approach, demonstrating its ability to generate high-fidelity 3D content without succumbing to issues such as over-smoothness.
@@ 1 Introduction
3D generative modelling is inherently more complex than 2D modelling, requiring meticulous attention to view-consistent fine geometry and texture. Despite the increased intricacy, 3D data is not as readily available as its 2D image counterparts, which have propelled recent advances in text-to-image generation [27, 29, 28]. Even with recent efforts of Objaverse [5], the uncurated 3D data only amounts to 10 million instances, in sharp contrast to the vast 5 billion image-text pairs available [30]. As a result, utilizing 2D supervision for 3D generation has emerged as a prominent research direction for text-to-3D generation. Notably, Score Distillation Sampling (SDS) [24, 38] is proposed to optimize a 3D representation so that its renderings at arbitrary views are highly plausible as evaluated by a pre-trained text-to-image model, without requiring any 3D data. As presented in DreamFusion [24], SDS enables the production of intriguing 3D models given arbitrary text prompts, but the results tend to be over-smooth and implausible (e.g., floaters). Subsequent works build upon SDS and enhance the generation quality with improvements in training practices. This is most effectively achieved by adopting higher resolution training, utilizing larger batch sizes, and implementing a coarse-to-fine generation approach with mesh optimization for sharper generation details [17, 3, 39]. While SDS remains fundamental and ubiquitously used for 3D generation, its theoretical understanding remains obscured and requires further exploration.
In this paper, we make extensive efforts to resolve the issues above, our contributions can be summarized as follows:
@@
  • We offer a comprehensive understanding of Score Distillation Sampling (SDS) by interpreting the noise residual term (referred to as the SDS estimator) as a combination of three functional components: mode-disengaging, mode-seeking and variance-reducing terms. Based on our analysis, we identify that the over-smoothness and implausibility problems in 3D generation arise from intrinsic deficiency of the first two terms. Moreover, we identify the main training signal to be the mode-disengaging term.
  • We show that the variance reduction scheme introduced by SDS incurs scale and direction mismatch problems, and propose a more advanced one.
@@
  • We propose Stable Score Distillation (SSD), which utilizes a simple yet effective term-balancing scheme that significantly mitigates the fundamental deficiencies of SDS, particularly alleviating issues such as over-smoothness and implausibility.
@@
  • We establish meaningful connections between our analytical findings and prevalent observations and practices in optimization-based 3D generation, such as the adoption of a large CFG scale and the generation of over-smoothed and floater-abundant results. These connections provide valuable perspectives for enhancing the optimization process of 3D generation.
@@
  • Extensive experiments demonstrate that our proposed method significantly outperforms baselines and our improvement is compatible with existing 3D generation frameworks and representations. Our approach is capable of high-quality 3D generation with vibrant colors and fine details.
Figure 1: 3D Gaussian generation from text prompts.
Figure 2: NeRF generation from text prompts.
Figure 3: Incorporating SSD to existing 3D generation frameworks consistently improves generation quality.
Figure 4: Comparisons between SSD and SOTA methods on text-to-3D generation. Baseline results are obtained from their papers.
Figure 5: More comparisons between SSD and SOTA methods on text-to-3D generation with more diverse prompts. For each text prompt, baseline results are obtained from theirs papers except for Fantasia3D, and presented on the left, while SSD results are shown on the right.
Figure 6: Ablation study on the design choices of SSD. Better viewed when zoomed in.
@@ 2 Related Works
Recently, DreamFusion [24] and SJC [38] propose to generate 3D content by optimizing a differentiable 3D representation so that its renderings at arbitrary viewpoints are deemed plausible by 2D diffusion priors. Such methods are commonly referred to as the optimization-based 3D generation. Subsequent works enhance the generation quality by utilizing 3D-aware diffusion models for image-to-3D generation [36, 26, 15, 20]; adopting a coarse-to-fine generation approach with improved engineering practices such as higher-resolution training and mesh optimization [17, 3]; exploring timestep scheduling [11]; or introducing additional generation priors [12, 1, 13, 16, 44]. ProlificDreamer [39], along with recent works of NFSD [14] and CSD [43] concurrent to ours provide new perspectives on SDS. however, there remains a gap in a comprehensive understanding of the SDS formulation, which is the focus of this work.
@@ 3.4 Score Distillation Sampling (SDS)
where $\boldsymbol{x}=g(\theta,c)$ and $w(t)$ re-scales the weights of gradients according to $t$, and $p_{\phi}(\boldsymbol{x}_{t};y,t)$ is the distribution of $\boldsymbol{x}_{t}$ implicitly represented by the score function $\hat{\boldsymbol{\epsilon}_{\phi}}$. DreamFusion [24] introduces the $\boldsymbol{\epsilon}$ term for variance reduction. We call the resultant noise residual term highlighted in red the SDS estimator for short.
@@ 4 Method
We start with problem setting in Sec. 4.1 and decompose the SDS estimator to three functional terms as shown in Eq. 4: the mode-disengaging term (denoted as $h$ for its homogeneous form), the mode-seeking term and the variance-reducing term. In Sec. 4.2, Sec. 4.3, and Sec. 4.4 we analyze the mathematical properties, numerical characteristics and intrinsic limitations of each term, respectively. Based on our analysis, we propose Stable Score Distillation (SSD) in Sec. 4.5 which efficiently utilizes the distinctive properties of each term to augment each other for higher-quality 3D content generation.
@@ 4.2.1 Mathematical Properties
Substituting Eq. 2 into the formulation of $h$, we obtain $h=-\sigma_{t}(\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)-\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};t))$. When optimizing a 3D representation $\theta$, the sub-terms of $h$, namely $-\sigma_{t}\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$ and $\sigma_{t}\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,\emptyset)$, respectively drives $\theta$ such that $\mathbb{E}_{\boldsymbol{\epsilon}}[\boldsymbol{x}_{t}]=\alpha_{t}\boldsymbol{x}$ will: a) approach the conditional modes of $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$, and b) distance the unconditional modes of $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};t)$ which are independent of the conditional prompt (for this reason we call $h$ mode-disengaging). We emphasize property b) to be critical, as it guides $\theta$ to avoid the over-smoothing problem caused by trapping points and transient modes, a concern encountered by the mode-seeking term (see Sec. 4.3 for definition and analysis). In Fig. 7 (left), we illustrate that the mode-disengaging term has the capability to navigate out of the trapping point of the Mode-Seeking term induced from its transient modes, thanks to the disengaging force that avoids unconditional modes. This capability is crucial for effective 3D generation. Considering that in Eq. 4, only the mode-disengaging and mode-seeking terms provide supervision signals (note that $\mathbb{E}_{\boldsymbol{\epsilon}}[\boldsymbol{\epsilon}]=\boldsymbol{0}$), our analysis positions $h$ as the main supervision signal in 3D generation. Our analysis provides theoretical explanation to a concurrent work, CSD [43], which also claims $h$ to be the main supervision signal.
@@ 4.2.2 Intrinsic Limitation
Based on previous discovery, one may be tempted to employ $h$ as the sole supervision signal in Eq. 3 for 3D generation, akin to CSD [43]. However, our experiments reveal a persistent issue of color-saturation in even 2D content generation when relying solely on $h$ (SSD, M=0 in Fig. 20). To delve deeper into this issue, we conduct experiments with varying values of $t$ and identify its particular association with small $t$, e.g., $t<200$ (see App. Fig. 15). Note that as $t\rightarrow 0$, $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)\rightarrow\mathbb{P}_{\phi}(\boldsymbol{x};y,t)$ and $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};\emptyset,t)\rightarrow\mathbb{P}_{\phi}(\boldsymbol{x};\emptyset,t)$. As the term $h$ seeks to maximize $\frac{\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)}{\mathbb{P}_{\phi}(\boldsymbol{x}_{t};\emptyset,t)}$, when $t\rightarrow 0$, the learning target of $h$ gradually approaches maximizing $\frac{\mathbb{P}_{\phi}(\boldsymbol{x};y)}{\mathbb{P}_{\phi}(\boldsymbol{x})}$. However, the maximizing points for this ratio are often not around any modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y)$, where $\mathbb{P}_{\phi}(\boldsymbol{x})$ is also high according to the mode consistency hypothesis. Worse still, they tend to be not around any singular modes of $\mathbb{P}_{\phi}(\boldsymbol{x})$, where $\mathbb{P}_{\phi}(\boldsymbol{x})$ is high and $\mathbb{P}_{\phi}(\boldsymbol{x};y)$ is low. In other words, when $t\rightarrow 0$, the mode-disengaging term discourages $g(\theta,c)$ from converging to any mode of a natural image distribution $\mathbb{P}_{\phi}(\boldsymbol{x})$, let alone to modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y)$, violating the necessary condition of 3D generation as presented in Sec. 4.1. Therefore, special treatments are necessary to enable the 3D representation to converge to modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y)$ when $t$ is small. We illustrate this problem in Fig. 7 (left), where the mode-disengaging trajectory converges to a point far from modes of $\mathbb{P}(\boldsymbol{x};y)$.
@@ 4.3.1 Numerical Properties
Scale. Recall $||\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)||=\sigma_{t}\cdot||\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)||$. Assuming $||\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)||$ is bounded, one should expect $||\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)||$ to increase as $t$ increases. We validate this assumption numerically in Appendix Fig. 16.
Direction. As $t\rightarrow T$, $\nabla_{\boldsymbol{x}_{t}}\text{log}\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$ gradually transforms from the score of conditional distribution of natural images, i.e., $\nabla_{\boldsymbol{x}}log\mathbb{P}_{\phi}(\boldsymbol{x};y)$, to $\nabla_{\boldsymbol{x}_{T}}\mathcal{N}(\boldsymbol{x}_{T};\boldsymbol{0},\boldsymbol{I})$, which are independent of $\boldsymbol{\epsilon}$ and collinear to $\boldsymbol{\epsilon}$ respectively (proved in Sec. 10). Therefore the linear correlation between $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)$ and $\boldsymbol{\epsilon}$ as defined in [25] is expected to increase as $t$ increases. We validate this assumption in Fig. 17.